OpenStudy (anonymous):

x3+2x-1 find complex roots

OpenStudy (solomonzelman):

there is a big retarded formula for \(\large\color{slate}{ ax^3+bx-c }\), I don't remember it and don't use it though.

OpenStudy (michele_laino):

the real root, is given by the subsequent formula: \[x=\sqrt[3]{q+\sqrt{q ^{2}-p ^{3}}}+\sqrt[3]{q-\sqrt{q ^{2}-p ^{3}}}\] wherep=-2/3, and q=1/2